# On the Age Calibration of Open Clusters using Red Clump Stars

### v2

@@ 2 Cluster Sample and Membership Analysis
(log years) (mag) (dex) (mag) (deg)
Melotte 71 8.99 ± $\pm\ 0.175$ 0.139 -0.22 0.118 ± $\pm\ 0.035$ 8 0.224
King 5 9.01 ± $\pm\ 0.175$ 0.897 -0.30 0.104 ± $\pm\ 0.011$ 8 0.272
NGC 2477 9.05 ± $\pm\ 0.175$ 0.390 -0.192 0.176 ± $\pm\ 0.008$ 58 0.450
NGC 1245 9.08 ± $\pm\ 0.175$ 0.335 0.10 0.107 ± $\pm\ 0.008$ 38 0.326
NGC 6208 9.15 ± $\pm\ 0.175$ 0.279 -0.03 0.082 ± $\pm\ 0.009$ 6 0.486
NGC 2509 9.18 ± $\pm\ 0.175$ 0.139 0.082 0.098 ± $\pm\ 0.014$ 3 0.179
NGC 7789 9.19 ± $\pm\ 0.175$ 0.296 -0.24 0.113 ± $\pm\ 0.005$ 35 0.422
NGC 2506 9.22 ± $\pm\ 0.175$ 0.056 -0.20 0.084 ± $\pm\ 0.004$ 33 0.263
NGC 2420 9.24 ± $\pm\ 0.175$ 0.013 -0.38 0.079 ± $\pm\ 0.004$ 7 0.210
Ruprecht 68 9.26 ± $\pm\ 0.15$ 0.446 0.13 0.151 ± $\pm\ 0.022$ 9 0.221
NGC 2627 9.27 ± $\pm\ 0.15$ 0.139 -0.12 0.066 ± $\pm\ 0.010$ 10 0.260
Melotte 66 9.63 ± $\pm\ 0.15$ 0.167 -0.33 0.123 ± $\pm\ 0.006$ 20 0.262
NGC 2682 9.63 ± $\pm\ 0.15$ 0.067 -0.102 0.103 ± $\pm\ 0.004$ 7 0.622
NGC 6791 9.80 ± $\pm\ 0.15$ 0.157 0.32 0.161 ± $\pm\ 0.005$ 36 0.204
@@
Name Age E($B-V$) [Fe/H] $d_{BP-RP}$ N(RC) Cluster radius
(log years) (mag) (dex) (mag) (deg)
SMC 26 0.03 0.160 ± $\pm\ 0.005$ 211 0.025
SMC 225 0.05 0.168 ± $\pm\ 0.011$ 29 0.004
SMC 245 0.03 0.122 ± $\pm\ 0.008$ 109 0.006
SMC 368 0.02 0.148 ± $\pm\ 0.009$ 65 0.007
SMC 571 0.08 0.145 ± $\pm\ 0.005$ 344 0.011
Average 8.85 ± $\pm\ 0.07$0.6 0.148 ± $\pm\ 0.004$
@@ 3 Results
An example of a cleaned CMD appears in Figure 5 and the remainder of cluster CMDs appear in Appendix A. From the CMDs we compute the central red clump color-magnitude location using the Tukey biweight statistic (Beers et al., 1990), which is similar to the median, in both $BP-RP$ and $G$. The error in central location is estimated as $\sigma=\sqrt{v/N}$, where $v$ is the biweight midvariance and $N$ is the number of red clump stars used for the calculation. RGB colors at the $G$ magnitude of the red clump are found by overlaying the RGB segment of a PARSEC (Bressan et al., 2012) isochrone on the data, using synthetic photometry from the Worthey models to generate the $BP-RP$ color. The synthetic RGB is then shifted left and right until fifty percent of stars were bluer and fifty percent of stars were redder. Once placed, the $BP-RP$ magnitude at the $G$ magnitude of the red clump is read from the shifted synthetic RGB. The clump stars outnumbered the RGB stars, so we expect counting statistics on the RGB to be the largest error source. We estimated RGB placement errors by shifting the theoretical RGB $\pm$ one star from the median position, and taking the average of the absolute value of those two shifts.
Unsolved astrophysical effects and the placement of the initial red clump selection box may slightly skew the Tukey biweight for finding the central location. An example of an astrophysical effect occurs for clusters where the initial mass of a star at the RGB tip is $\sim 2$ M, of which our sample contains one example, NGC 7789. This cluster and NGC 752 (Girardi et al., 2000) have a clump with a few extra stars that are both lower luminosity and slightly blue. Significant spread in mass loss on the RGB combined with the fact that this mass might straddle the mass at which helium burning begins in a degenerate versus a nondegenerate core might explain the effect. An example at older ages but high metallicity is NGC 6791, some of whose RGB stars may skip helium burning altogether (Hansen, 2005).
@@
Figure 7: Color delta versus log age for galactic open clusters (circles) and average of SMC fields (right-facing triangles) from this work, along with globular clusters (squares), SMC clusters (thin diamonds), and LMC clusters (wide diamonds) from H91 color mapped by metallicity as indicated by the color bar. The H91 linear fit (dashed line) is also shown.
@@ 4 Discussion and Conclusion
Even restricted to the new measurements, and further restricted to log age $<9.5$, the scatter appears to be astrophysical. A Kolmogorov-Smirnov test on our $d_{BP-RP}$ values for clusters in our sample younger than 9.5 log years returned $p\approx 1.62\times 10^{-5}$, indicating that the scatter is not drawn from a normal distribution. The cause of the scatter cannot be age, or we would see a slope in Figure 7. We sample metallicity span less well, but within that caveat, no metallicity dependence is evident. If neither age nor metallicity is the cause of the modulation in $d_{B-R}$, we are driven to consider more subtle causes such as variation in helium abundance, variation in the [$\alpha$/Fe] ratio, or variation in binary separation distributions. The investigation into the cause of the scatter must await future work.
@@
We examine some predictions from stellar evolution theory in Figure 8. All isochrones were translated from $T_{\rm eff}$ and log $L$ to $B-R$ using the empirical transformations of Worthey & Lee (2011). We show BaSTI 5.0.1 (Pietrinferni et al., 2004), BaSTI updated (Hidalgo et al., 2018), Padova, retrieved 2011 (Marigo et al., 2008), at two abundance values, and PARSEC+COLIBRI (Bressan et al., 2012; Pastorelli et al., 2020) with Riemers (Reimers, 1977) $\eta=0.2$, retrieved 2023, also at two abundance values. Within the isochrone files, red clumps were located by treating the isochrone as a stellar population, predicting relative numbers of stars, and choosing the dominant sampling point to correspond to the red clump location. On the positive side, all evolutionary sets are within $\sim$0.1 mag (about 150 K) of the observed color delta, and all sets show a growth in $d_{B-R}$ for ancient stellar populations that qualitatively mimics the observations.
If precision better than 150 K are desired, theory does not match observation. In particular, reading from Figure 8, theory predicts a metallicity dependence of roughly $\frac{d(B-R)}{d{\rm[Fe/H]}}\geq-0.2$ mag per decade in heavy element abundance. A metallicity dependence this large is strictly ruled out by the data, which includes clusters from the large and small Magellanic Clouds. For example, super-metal-rich cluster NGC 6791 lies right among SMC clusters. The presence of this systematic among current isochrone sets affects CMD decompositions (Harris & Zaritsky, 2001; Dolphin, 2002). It also impacts integrated light models by introducing a fictitious color drift in the sense to make metal rich populations redder.
@@
Modelers should be aware of this discrepancy between current stellar evolution theory and observation because it adds to a list of ill-modeled or un-modeled effects that are far larger than the RMS fit between galaxy spectra and model integrated light spectra. Conroy et al. (2014), for example, find fits as good as $0.2\%$ RMS, but their red clump temperatures do not adhere to H91 and so could be off by several percent. Blue stragglers are not included, which can lead to errors of over a magnitude in the terrestrial UV (Deng et al., 1999). Chemically peculiar stars, not included, alter the blue spectrum by $\sim 2\%$ in the blue (Worthey & Shi, 2023). Binary evolution products, carbon stars, and metallicity-composite populations are likewise not modeled, and yet the optical spectrum matches to $0.2\%$. The reason the fit is excellent is a generalization the age-metallicity degeneracy (Worthey, 1994), where instead of age or metallicity, one inserts the concept of stellar temperatures in general. While some astrophysical inferences, such as abundance ratios, are somewhat isolated from this degeneracy, others, such as age, metallicity, age spread, or metallicity spread are not.
@@ Appendix A Color Magnitude Diagrams
The CMDs of all fourteen Milky Way open clusters and five SMC fields used in this work are presented here. For each figure, we zoom in on the RGB and include the estimated boundary of the red clump, the fitted RGB segment of a solar-metallicity PARSEC isochrone, and the red clump and RGB loci to illustrate the method.

### v3

@@ 2 Cluster Sample and Membership Analysis
(log years) (mag) (dex) (mag) (mag) (deg)
Melotte 71 8.99 ± $\pm\ 0.175$ 0.139 -0.22 0.118 ± $\pm\ 0.035$ 0.208 ± $\pm\ 0.046$ 8 0.224
King 5 9.01 ± $\pm\ 0.175$ 0.897 -0.30 0.104 ± $\pm\ 0.012$ 0.185 ± $\pm\ 0.016$ 8 0.272
NGC 2477 9.05 ± $\pm\ 0.175$ 0.390 -0.192 0.176 ± $\pm\ 0.009$ 0.311 ± $\pm\ 0.013$ 58 0.450
NGC 1245 9.08 ± $\pm\ 0.175$ 0.335 0.10 0.107 ± $\pm\ 0.009$ 0.189 ± $\pm\ 0.012$ 38 0.326
NGC 6208 9.15 ± $\pm\ 0.175$ 0.279 -0.03 0.082 ± $\pm\ 0.010$ 0.146 ± $\pm\ 0.014$ 6 0.486
NGC 2509 9.18 ± $\pm\ 0.175$ 0.139 0.082 0.098 ± $\pm\ 0.015$ 0.175 ± $\pm\ 0.020$ 3 0.179
NGC 7789 9.19 ± $\pm\ 0.175$ 0.296 -0.24 0.113 ± $\pm\ 0.006$ 0.195 ± $\pm\ 0.010$ 35 0.422
NGC 2506 9.22 ± $\pm\ 0.175$ 0.056 -0.20 0.084 ± $\pm\ 0.006$ 0.149 ± $\pm\ 0.009$ 33 0.263
NGC 2420 9.24 ± $\pm\ 0.175$ 0.013 -0.38 0.079 ± $\pm\ 0.006$ 0.14 ± $\pm\ 0.009$ 7 0.210
Ruprecht 68 9.26 ± $\pm\ 0.15$ 0.446 0.13 0.151 ± $\pm\ 0.022$ 0.240 ± $\pm\ 0.030$ 9 0.221
NGC 2627 9.27 ± $\pm\ 0.15$ 0.139 -0.12 0.066 ± $\pm\ 0.011$ 0.117 ± $\pm\ 0.015$ 10 0.260
Melotte 66 9.63 ± $\pm\ 0.15$ 0.167 -0.33 0.123 ± $\pm\ 0.007$ 0.217 ± $\pm\ 0.011$ 20 0.262
NGC 2682 9.63 ± $\pm\ 0.15$ 0.067 -0.102 0.103 ± $\pm\ 0.006$ 0.170 ±$\pm 0.009$ 7 0.622
NGC 6791 9.80 ± $\pm\ 0.15$ 0.157 0.32 0.161 ± $\pm\ 0.006$ 0.183 ± $\pm\ 0.009$ 36 0.204
@@
Name Age E($B-V$) [Fe/H] $d_{BP-RP}$ $d_{B-R}$ N(RC) Cluster radius
(log years) (mag) (dex) (mag) (mag) (deg)
SMC 26 0.03 0.120 ± $\pm\ 0.007$ 0.212 ± $\pm\ 0.010$ 211 0.025
SMC 225 0.05 0.125 ± $\pm\ 0.010$ 0.222 ± $\pm\ 0.016$ 29 0.004
SMC 245 0.03 0.091 ± $\pm\ 0.008$ 0.161 ± $\pm\ 0.013$ 109 0.006
SMC 368 0.02 0.110 ± $\pm\ 0.009$ 0.195 ± $\pm\ 0.014$ 65 0.007
SMC 571 0.08 0.109 ± $\pm\ 0.007$ 0.192 ± $\pm\ 0.010$ 344 0.011
Average 8.85 ± $\pm\ 0.07$0.6 0.111 ± $\pm\ 0.004$ 0.196 ± $\pm\ 0.007$
@@ 3 Results
An example of a cleaned CMD appears in Figure 5 and the remainder of cluster CMDs appear in Appendix A. From the CMDs we drew boxes by hand and computed the central red clump color-magnitude location using the Tukey biweight statistic (Beers et al., 1990), which is similar to the median, in both $BP-RP$ and $G$. The error in central location is estimated as $\sigma=\sqrt{v/N}$, where $v$ is the biweight midvariance and $N$ is the number of red clump stars used for the calculation. RGB colors at the $G$ magnitude of the red clump are found by overlaying the RGB segment of a PARSEC (Bressan et al., 2012) isochrone on the data, using synthetic photometry from the Worthey models to generate the $BP-RP$ color. The synthetic RGB is then shifted left and right until fifty percent of stars were bluer and fifty percent of stars were redder. Once placed, the $BP-RP$ magnitude at the $G$ magnitude of the red clump is read from the shifted synthetic RGB. The clump stars outnumbered the RGB stars, so we expect counting statistics on the RGB to be the largest error source. We estimated RGB placement errors by shifting the theoretical RGB $\pm$ one star from the median position, and taking the average of the absolute value of those two shifts. For small $N$, our biggest concern, this is a good error estimate assuming the nearest neighbor distribution function from a Poisson point process. For large $N$, the method converges faster than $N^{-1/2}$ and become inappropriate, but there is also a theoretical error floor from differences in RGB tilts between different isochrone sets. We compared isochrones of 1 Gyr age and solar composition from Worthey (1994); Pietrinferni et al. (2004); Hidalgo et al. (2018); Bressan et al. (1994); Marigo et al. (2008); Bressan et al. (2012) and Dotter et al. (2007) and performed statistics on the RGB slopes to arrive at rough error floors of 0.004 mag for $d(BP-RP)$ and 0.007 mag for $d(B-R)$. These errors were added in quadrature with the nearest neighbor errors for the final value tabulated in Tables 1 and 2.
Unsolved astrophysical effects and the placement of the initial red clump selection box may slightly skew the Tukey biweight for finding the central location. An example of an astrophysical effect occurs for clusters where the initial mass of a star at the RGB tip is $\sim 2$ M. NGC 7789 and NGC 752 (which is not in our sample) (Girardi et al., 2000) have a clump with a few extra stars that are both lower luminosity and slightly blue. Significant spread in mass loss on the RGB combined with the fact that this mass might straddle the mass at which helium burning begins in a degenerate versus a nondegenerate core might explain the effect. Many of our clusters have turnoff masses near 2 M, and some (NGC 1245, NGC 2477) have “tall” red clumps not predicted by theory that could potentially be explained by this effect. On the other hand, theory is currently unable to account for fundamental red clump properties (degl’Innocenti et al., 2000; An et al., 2019). The mass loss history, be it pre-flash or flash-caused, is frequently invoked as a possible explanation. A separate example at older ages but high metallicity is NGC 6791, some of whose RGB stars may skip helium burning altogether (Hansen, 2005).
A “tall” clump also generates a new error source, namely uncertainty in the luminosity of clump. Transferred to the RGB, the luminosity uncertainty translates to a color uncertainty, albeit a mild one due to the steepness of the RGB. This error was included by computing a biweight location and scale in the $G$ dimension and scaling it appropriately via the $dG/d(BP-RP)$ RGB slope.
@@
Figure 7: Color delta versus log age for galactic open clusters (circles) and average of SMC fields (right-facing triangles) from this work, along with galactic open clusters (squares), galactic globular clusters (pointing-up triangles), SMC clusters (thin diamonds), and LMC clusters (wide diamonds) from H91 color mapped by metallicity as indicated by the color bar. The H91 linear fit (dashed line) is also shown.
@@ 4 Discussion and Conclusion
Even restricted to the new measurements, and further restricted to log age $<9.5$, the scatter appears to be astrophysical. A Kolmogorov-Smirnov test on our $d_{BP-RP}$ values for clusters in our sample younger than 9.5 log years returned $p\approx 0.002$, indicating that the scatter is not drawn from a normal distribution. The cause of the scatter cannot be age, or we would see a slope in Figure 7. We sample metallicity span less well, but within that caveat, no metallicity dependence is evident. If neither age nor metallicity is the cause of the modulation in $d_{B-R}$, we are driven to consider more subtle causes such as variation in helium abundance, variation in the [$\alpha$/Fe] ratio, or variation in binary separation distributions. The investigation into the cause of the scatter must await future work.
@@
We examine some predictions from stellar evolution theory in Figure 8. All isochrones were translated from $T_{\rm eff}$ and log $L$ to $B-R$ using the empirical transformations of Worthey & Lee (2011). We show BaSTI 5.0.1 (Pietrinferni et al., 2004), BaSTI updated (Hidalgo et al., 2018), Padova, retrieved 2011 (Marigo et al., 2008), at two abundance values, and PARSEC+COLIBRI (Bressan et al., 2012; Pastorelli et al., 2020) with Riemers (Reimers, 1975) $\eta=0.2$, retrieved 2023, also at two abundance values. Within the isochrone files, red clumps were located by treating the isochrone as a stellar population, predicting relative numbers of stars, and choosing the dominant sampling point to correspond to the red clump location. On the positive side, all evolutionary sets are within $\sim$0.1 mag (about 150 K) of the observed color delta, and all sets show a growth in $d_{B-R}$ for ancient stellar populations that qualitatively mimics the observations.
If precision better than 150 K is desired, theory does not match observation. In particular, reading from Figure 8, theory predicts a metallicity dependence of roughly $\frac{d(B-R)}{d{\rm[Fe/H]}}\geq-0.2$ mag per decade in heavy element abundance. A metallicity dependence this large is strictly ruled out by the data, which includes clusters from the large and small Magellanic Clouds. For example, super-metal-rich cluster NGC 6791 lies right among SMC clusters. The presence of this systematic among current isochrone sets affects CMD decompositions (Harris & Zaritsky, 2001; Dolphin, 2002). It also impacts integrated light models by introducing a fictitious color drift in the sense to make metal rich populations redder.
@@
Modelers should be aware of this discrepancy between current stellar evolution theory and observation because it adds to a list of ill-modeled or un-modeled effects that are far larger than the RMS fit between galaxy spectra and modeled integrated light spectra. Conroy et al. (2014), for example, find fits as good as $0.2\%$ RMS, but their red clump temperatures do not adhere to H91 and so could be off by several percent. Blue stragglers are not included, which can lead to errors of over a magnitude in the terrestrial UV (Deng et al., 1999). Chemically peculiar stars, not included, alter the blue spectrum by $\sim 2\%$ in the blue (Worthey & Shi, 2023). Binary evolution products, carbon stars, and metallicity-composite populations are likewise not modeled, and yet the optical spectrum matches to $0.2\%$. The reason the fit is excellent is a generalization the age-metallicity degeneracy (Worthey, 1994), where instead of age or metallicity, one inserts the concept of stellar temperatures in general. While some astrophysical inferences, such as abundance ratios, are somewhat isolated from this degeneracy, others, such as age, metallicity, age spread, or metallicity spread are not.
@@ Appendix A Color Magnitude Diagrams
The CMDs of all fourteen Milky Way open clusters and five SMC fields used in this work are presented in Figs 10, 11, and 12. In each figure, we zoom in on the RGB and include the estimated boundary of the red clump, the fitted RGB segment of a solar-metallicity PARSEC isochrone, and the red clump and RGB loci to illustrate the method.